OpenStudy (anonymous):

The concentration of glucose in the aqueous fluid of the spine is 75.0 mg/100mL of fluid. What is the molar concentration of glucose in spinal fluid?

OpenStudy (frostbite):

\[\huge \rho = \frac{ m }{ V }=\frac{ M \times n }{ V }=C \times M\] Can you solve it from here?

OpenStudy (anonymous):

Trying it now

OpenStudy (anonymous):

0.075g/180.16g/mol = 4.16x10^-4 mol glucose 100mL = 0.1 L 4.16x10^-4 mol / 0.1 L = 4.16x10^-3 M is this right?

OpenStudy (frostbite):

Looks correct to me.

OpenStudy (anonymous):

Ok thank you

OpenStudy (frostbite):

No problem at all.